# Aeos: Transport of metals from minihalos following Population III stellar feedback

### v1

@@ 2.2.2 Population II Star Formation
Pop II star formation occurs when the metallicity threshold described for Pop III stars is exceeded, and does not require an H2 threshold. Pop II stellar masses are sampled from a Kroupa IMF (Kroupa, 2001) with a minimum mass of 0.08 $\rm M_{\odot}$ and a maximum mass of 120 $\rm M_{\odot}$. Stars above 2 $\rm M_{\odot}$ are treated as individual star particles, however, stars below that threshold in a star forming cell are grouped together into a single particle as they do not contribute to feedback or metal enrichment on the timescales of this simulation. We set stellar radii, effective temperature, surface gravity, lifetime, and AGB phase using the stellar evolution data from PARSEC (Bressan et al., 2012).
@@ 2.3 Stellar Feedback
Stellar Winds: There are two types of stellar winds: asymptotic giant branch (AGB) stellar winds and winds from massive stars. Winds are deposited onto the grid over a 2 pc radius. As resolving the fast hot winds is computationally expensive (see E19), the wind velocity for all massive stars is fixed over their lifetimes at a maximum value of 100 km s-1. The kinetic energy of the winds is fully thermalized before injection. Mass loss rates are assumed to be constant for all winds over the lifetime of the star.
Supernovae: We implement two different types of supernovae: CCSNe and SNIa. With a maximum resolution of 1 proper pc in these simulations, we are able to resolve the Sedov-Taylor phase of the vast majority of our supernovae at the typical gas densities in which they explode (Smith et al., 2018; Hu, 2019). Therefore, for these events, we only include the thermal energy deposition. Energy and mass from each supernova explosion are deposited over a 2 pc radius sphere.
Core-Collapse Supernovae: Pop III stars of masses $10\,\rm M_{\odot}$ $<M_{*}\,<$ $100\,\rm M_{\odot}$ and Pop II stars of masses $8\,\rm M_{\odot}$ $<\,M_{*}\,<$ $25\,\rm M_{\odot}$ explode as CCSNe with a fixed energy of $10^{51}$ erg. Stars above these ranges undergo direct collapse to a black hole at the end of their lifetimes with no mass or energy feedback. Though pair-instability supernovae are expected to occur for stars in the range of $140-260\rm M_{\odot}$, we do not include these in our current simulations.
Type Ia Supernovae: Pop II stars $M_{*}<8\,\rm M_{\odot}$ become white dwarfs, with stars $3\,{\rm M_{\odot}}<M_{*}<8\,\rm M_{\odot}$ being capable of exploding as SNIa. For SNIa, we use the "standard" delay time distribution model in Ruiter et al. (2011), which sums four different channels: 1) double degenerate, 2) single degenerate, 3) helium-rich donor, and 4) a sub-Chandrasekhar mass. We use a random number draw from this distribution to determine when and if a white dwarf will explode as a SNIa. SNIa explode with a fixed energy of $10^{51}$ erg.
@@ 2.4 Stellar Yields
We capture detailed chemical yields for Pop III and Pop II CCSNe, SNIa, and stellar winds including C, N, O, Na, Mg, Ca, Mn, Fe, Sr, and Ba metal abundances in addition to H, He, and overall metallicity, Z. The individual metal abundances trace different channels of feedback and progenitor mass. O, Mg, and Ca primarily trace CCSNe, and trace short timescales of chemical evolution (10 Myr) (Timmes et al., 1995; Kobayashi et al., 2006). The evolution between SNIa and CCSNe is traced by the relative abundances of O, Mg, and Ca to Fe, which tracks chemical evolution on timescales of 100 Myr to 1 Gyr (Tinsley, 1980; Matteucci & Greggio, 1986; Hayden et al., 2015). S-process enrichment from low-mass AGB stars is traced using N, Ba (4-8 $\rm M_{\odot}$), and Sr (<4 $\rm M_{\odot}$) (e.g. Karakas & Lattanzio, 2014, and references therein).
@@
Pop III CCSNe Yields: Pop III CCSNe ($10\,\rm M_{\odot}$ $<\,M_{*}\,<$ $100\,\rm M_{\odot}$) yields are adopted from Heger & Woosley (2010) using standard 0.1 mixing. It is reasonable to approximate that all stars up to 100 $\rm M_{\odot}$ end their lives as CCSNe, though stars in the range 70-120 $\rm M_{\odot}$ may end their lives as either CCSNe or by directly collapsing to a black hole (Woosley, 2017).
@@
Pop II CCSNe Yields: Yields for Pop II CCSNe were adopted from Limongi & Chieffi (2018) using 9 grid points in mass ($13\,\rm M_{\odot}$ $\leq\,M_{*}\,\leq$ $120\,\rm M_{\odot}$) and 4 grid points in metallicity ($3.236\times 10^{-5}\leq Z_{*}\leq 0.01345$). We use linear interpolation between grid points on the table. Stars with masses outside of the grid points adopt the abundance ratios of the nearest grid point and scale the yields linearly with mass. Stars with metallicities outside of the grid points adopt the yields of the closest grid point.
@@
Figure 3: Top: Fraction of injected metal mass that remains in halos where the boundary of the halo is either the virial radius (purple) or the contour where gas is denser than $\rho_{200/3}$, a density that is typical of collapsed halos (black). Center: The total metallicities (${M_{\rm met}/M_{\rm gas}}$) of halos (solid) and the IGM (dashed) using the $\rho_{200/3}$ density cut. Bottom: The average metallicity of hot (red), warm (orange), and cool (blue) gas at any density as well as star forming (black) gas. Though star formation begins much earlier in the simulation, star forming gas does not appear in this plot until 200 Myrs as the data is sparsely sampled in time and star formation is efficient. The metallicity of the star forming gas is expected to be no higher than that of cold gas at early times.
@@
To account for the yields from different stellar rotation models in Limongi & Chieffi (2018), we adopt a mixture model representing the population-averaged yields using the stellar rotation population fractions in Prantzos et al. (2018). Lastly, while these yields reasonably agree with stellar abundance observations in [X/Fe] vs. [Fe/H], Mg is under-produced, so we artificially increase the Mg yield from all massive stars by a factor of 2.2 (See Appendix of B24).
@@
Stellar Wind Yields: Yields for AGB winds were adopted from Cristallo et al. (2015) using 8 grid points in mass ($1.3\,\rm M_{\odot}$ $\leq\,M_{*}\,\leq$ $6.0\,\rm M_{\odot}$) and 10 grid points in metallicity ($1\times 10^{-4}\leq Z_{*}\leq 0.02$). The yields for massive stellar winds were also adopted from Limongi & Chieffi (2018) using the same grid points and interpolation as Pop II core-collapse supernovae (CCSNe).
@@
SNIa Yields: As discussed in Section 2.3, the delay time distribution of SNIa is calculated from a combination of four different SNIa sources. We assume a single abundance pattern for all SNIa (Thielemann et al., 1986), which in post-processing can be rescaled based on the number and type of SNIa that have occurred in the simulation. As a caveat, this assumes that the yields for each SNIa type do not affect the dynamical evolution of our box.
@@ 3 Metals Throughout Aeos
[…]
@@
To start, we consider the metallicity and fraction of gas phase metals in halos versus what is available the the IGM. To define the boundary of the halo-IGM interface, we choose to use a density cut at $\rho_{200/3}\,(\equiv\rho_{200}/3)$, as this approximates the density of gas in a virialized region in the absence of feedback (Navarro et al., 1997). Figure 2 shows examples of gas density and metal density slices of the most massive halos ($M>10^{7}\,\rm{M_{\odot}}$), with both the virial radius (given by Equation 1 below) and $\rho_{200/3}$ contour marked. Note that the $\rho_{200/3}$ density contour generally lies well within the virial radius for these halos. From these slices alone, it becomes apparent that a significant fraction of metals lie beyond the boundaries of halos, regardless of how they are defined.
@@
The top panel of Figure 3 shows the fraction of metal mass in halos $f_{\rm met}$, defining the boundary as either the virial radius or the $\rho_{200/3}$ density contour. Up until roughly 260 Myr ($z\approx 15.3$), $f_{\rm met}=0$–5% within the $\rho_{200/3}$ contour, or $f_{\rm met}\approx 8$% on average within the virial radius, emphasizing that most of the metals produced are expelled out of the halo, and more importantly, out of star forming regions. It is not until after 260 Myr that metals begin to reaccrete to, or are retained within, the most massive halos following SN explosions. Even then, the metal fraction only reaches $f_{\rm met}\sim$31–35%, however, this increase is rapid and may reach even higher fractions in a short amount of time following the end of the simulation. Of the metals in all halos, 92% reside in the four most massive halos in the simulation.
@@ 4 Metal Loss from Halos
The explosions of the first SNe quickly expel gas and metals from halos too small to retain their material. The fraction of metals retained by, or conversely, lost from each halo and the efficiency of mixing of the remaining metals determines when and how quickly star formation from enriched gas takes place in that halo or elsewhere.
[…]
@@
The metal loss fraction of each individual parent halo (halos that are not part of a more massive halo) relative to the halo mass is shown in Figure 2, color-coded by the number of SN that have exploded in the halo by the final redshift, $z=14.3$. Most halos have lost $\sim$100of their injected metals, which appears to be independent of the number of SNe that had previously exploded in that halo. Those halos that have not lost all metals can be divided into two groups:
@@ 4.1.1 Characteristic Evolution
Figure 6 shows the time evolution in dark matter, baryon fraction $f_{b}$, and metal mass respectively within the virial radius of the most massive progenitor of these four halos. Black circles represent instances of new SNe, with sizes scaled based on the number of new SNe. Magenta points indicate merger events.
All four halos show a similar dark matter mass evolution, originating very early on, within 50 Myr of each other and maintaining masses within one order of magnitude. Despite their similar dark matter evolution, they have widely varying gas evolution, with baryon fractions at times differing by up to five orders of magnitude. Rapid early decreases in ${f_{b}}$ are associated with the explosion of the first SNe.
Halos 1 and 2, which reach the lowest baryon fractions, have among the lowest dark matter masses at the minimum baryon fraction, hinting that the retention of material is tied to the depth of the potential well of the halo which is closely tied to $T_{\rm vir}$ and is redshift dependent. At their $f_{b}$ minima, Halos 1 and 2 have $T_{\rm vir}\approx 798$ K ($v_{c}\sim 4.7\,\rm km\,s^{-1}$) and $1056$ K ($v_{c}\sim 5.4\,\rm km\,s^{-1}$) respectively, indicating shallower potentials than Halos 0 and 3 which have $T_{\rm vir}\approx 3295$ ($v_{c}\sim 9.5\,\rm km\,s^{-1}$) and $2800$ K ($v_{c}\sim 8.8\,\rm km\,s^{-1}$) respectively. Though Halo 0 has a similar dark matter mass to the other halos at the time of its first SN, it is in a stage of rapid dark matter growth, which slows the loss of gas and brings in new gas, keeping the baryon fraction from dropping more than an order of magnitude.
The time evolution of metal mass also demonstrates interestingly similar characteristics between these four halos. All halos, at the time of their first SNe, show a rapid spike in their metal mass, immediately followed by a metal loss of 2–3 orders of magnitude. Similar characteristics can be seen in the metallicity evolution. Subsequent SNe over the next $\sim$ 25 Myr also result in the loss of metal mass, after which halos begin to steadily acquire metals and gas from accretion, SNe, and winds. Most importantly, SN explosions after this point no longer appear to drive metal mass out of the halo, although they also no longer appear to contribute greatly to the total metal mass of the halo. This may indicate that while the bulk of the metal mass, which is being accreted, is not being expelled from the halo, SNe yields are still escaping the halo. This is supported by the flat (or in the case of Halo 2, decreasing) metallicity with time, indicating that the ratios of metals to gas are not increasing as would be expected if yields injected by SNe were fully retained. As a result, metallicities only begin to increase near the end of the simulation.

### v2

@@ 2.2.2 Population II Star Formation
Pop II star formation occurs when the metallicity threshold described for Pop III stars is exceeded, and does not require an H2 threshold. Pop II stellar masses are sampled from a Kroupa IMF (Kroupa, 2001) with a minimum mass of 0.08 $\rm M_{\odot}$ and a maximum mass of 120 $\rm M_{\odot}$. Stars above 2 $\rm M_{\odot}$ are treated as individual star particles; however, stars below that threshold in a star forming cell are grouped together into a single particle as they do not contribute to feedback or metal enrichment on the timescales of this simulation. We set stellar radii, effective temperature, surface gravity, lifetime, and asymptotic giant branch (AGB) phase using the stellar evolution data from PARSEC (Bressan et al., 2012).
@@ 2.3 Stellar Feedback
Stellar Winds: There are two types of stellar winds: AGB stellar winds and winds from massive stars. Winds are deposited onto the grid over a 2 pc radius. As resolving the fast hot winds is computationally expensive (see E19), the wind velocity for all massive stars is fixed over their lifetimes at a maximum value of 100 km s-1. The kinetic energy of the winds is fully thermalized before injection. Mass loss rates are assumed to be constant for all winds over the lifetime of the star.
Supernovae: We implement two different types of SNe: CCSNe and SNIa. With a maximum resolution of 1 proper pc in these simulations, we are able to resolve the Sedov–Taylor phase of the vast majority of our SNe at the typical gas densities in which they explode (Smith et al., 2018; Hu, 2019). Therefore, for these events, we only include the thermal energy deposition. Energy and mass from each SN explosion are deposited over a 2 pc radius sphere.
Core-Collapse Supernovae: Pop III stars of masses $10\,\rm M_{\odot}$ $M_{*}\,<$ $100\,\rm M_{\odot}$ and Pop II stars of masses $8\,\rm M_{\odot}$ $M_{*}\,<$ $25\,\rm M_{\odot}$ explode as CCSNe with a fixed energy of $10^{51}$ erg. Stars above these ranges undergo direct collapse to a black hole at the end of their lifetimes with no mass or energy feedback. Although pair-instability SNe are expected to occur for stars in the range of $140-260\rm M_{\odot}$, we do not include these in our current simulations due to their rarity and unconfirmed existence.
Type Ia Supernovae: Pop II stars $M_{*}<8\,\rm M_{\odot}$ become white dwarfs, with stars $3\,{\rm M_{\odot}}<M_{*}<8\,\rm M_{\odot}$ being capable of exploding as SNIa. For SNIa, we use the "standard" delay time distribution model in Ruiter et al. (2011), which sums four different channels: (1) double degenerate, (2) single degenerate, (3) helium-rich donor, and (4) a sub-Chandrasekhar mass. We use a random number drawn from this distribution to determine when and if a white dwarf will explode as an SNIa. SNIa explode with a fixed energy of $10^{51}$ erg.
@@ 2.4 Stellar Yields
We capture detailed chemical yields for Pop III and Pop II CCSNe, SNIa, and stellar winds including C, N, O, Na, Mg, Ca, Mn, Fe, Sr, and Ba metal abundances in addition to H, He, and overall metallicity, Z. The individual metal abundances trace different channels of feedback and progenitor mass. O, Mg, and Ca primarily trace CCSNe, and trace short timescales of chemical evolution (10 Myr; Timmes et al., 1995; Kobayashi et al., 2006). The evolution between SNIa and CCSNe is traced by the relative abundances of O, Mg, and Ca to Fe, which tracks chemical evolution on timescales of 100 Myr to 1 Gyr (Tinsley, 1980; Matteucci & Greggio, 1986; Hayden et al., 2015). S-process enrichment from low-mass AGB stars is traced using N, Ba (4-8 $\rm M_{\odot}$), and Sr (<4 $\rm M_{\odot}$) (e.g. Karakas & Lattanzio, 2014, and references therein).
@@
Pop III CCSNe Yields: Pop III CCSNe ($10\,\rm M_{\odot}$ $M_{*}\,<$ $100\,\rm M_{\odot}$) yields are adopted from Heger & Woosley (2010) using standard 0.1 mixing. It is reasonable to approximate that all stars up to 100 $\rm M_{\odot}$ end their lives as CCSNe, although stars in the range 70-120 $\rm M_{\odot}$ may end their lives either as CCSNe or by directly collapsing to a black hole (Woosley, 2017).
@@
Pop II CCSNe Yields: Yields for Pop II CCSNe were adopted from Limongi & Chieffi (2018) using nine grid points in mass ($13\,\rm M_{\odot}$ $\leq\,M_{*}\,\leq$ $120\,\rm M_{\odot}$) and four grid points in metallicity ($3.236\times 10^{-5}\leq Z_{*}\leq 0.01345$). We use linear interpolation between grid points on the table. Stars with masses outside of the grid points adopt the abundance ratios of the nearest grid point and scale the yields linearly with mass. Stars with metallicities outside of the grid points adopt the yields of the closest grid point.
@@
Figure 3: Top: Fraction of injected metal mass that remains in halos where the boundary of the halo is either the virial radius (purple) or the contour where gas is denser than $\rho_{200/3}$, a density that is typical of collapsed halos (black). Center: The total metallicities ($M_{\rm met}/M_{\rm gas}$}) of halos (solid) and the IGM (dashed) using the $\rho_{200/3}$ density cut. Bottom: The average metallicity of hot (red), warm (orange), and cool (blue) gas at any density as well as star forming (black) gas. Alhough star formation begins much earlier in the simulation, star forming gas does not appear in this plot until 200 Myrs as the data are sparsely sampled in time, and star formation is efficient. The metallicity of the star forming gas is expected to be no higher than that of cold gas at early times.
@@
To account for the yields from different stellar rotation models in Limongi & Chieffi (2018), we adopt a mixture model representing the population-averaged yields using the stellar rotation population fractions in Prantzos et al. (2018). Lastly, while these yields reasonably agree with stellar abundance observations in [X/Fe] vs. [Fe/H], Mg is underproduced, so we artificially increase the Mg yield from all massive stars by a factor of 2.2 (see Appendix of B24).
@@
Stellar Wind Yields: Yields for AGB winds were adopted from Cristallo et al. (2015) using 8 grid points in mass ($1.3\,\rm M_{\odot}$ $\leq\,M_{*}\,\leq$ $6.0\,\rm M_{\odot}$) and 10 grid points in metallicity ($1\times 10^{-4}\leq Z_{*}\leq 0.02$). The yields for massive stellar winds were also adopted from Limongi & Chieffi (2018) using the same grid points and interpolation as Pop II CCSNe.
@@
SNIa Yields: As discussed in Section 2.3, the delay time distribution of SNIa is calculated from a combination of four different SNIa sources. We assume a single abundance pattern for all SNIa (Thielemann et al., 1986), which in postprocessing can be rescaled based on the number and type of SNIa that have occurred in the simulation. As a caveat, this assumes that the yields for each SNIa type do not affect the dynamical evolution of our box.
@@ 3 Metals throughout Aeos
Figure 5: Metal loss fraction, $\ell$. Top: Metal loss fraction at the last simulation snapshot at $z=14.3$ colored by the total number of SNe exploded in each halo. A metal loss fraction of $\ell=0$ indicates that the current metal mass $M_{\rm met}$ is equivalent to the total metal mass injected into the halo $M_{\rm met,inj}$. Bottom: Gas mass of halos, colored by fractional metal loss at $z=14.3$. Halos that have not had SNe (black points) generally have gas masses $M_{g}\sim f_{\rm bc}M_{\rm dm}$, where $f_{\rm bc}$ is the cosmological baryon fraction, $0.154$. Halos that have had SNe (larger colored points) have distinctly lower gas masses for the same dark matter masses $M_{\rm dm}$.
@@
To start, we consider the metallicity and fraction of gas phase metals in halos versus what is available in the IGM. To define the boundary of the halo–IGM interface, we choose to use a density cut at $\rho_{200/3}\,(\equiv\rho_{200}/3)$, as this approximates the density of gas in a virialized region in the absence of feedback (Navarro et al., 1997). Figure 2 shows examples of gas density and metal density slices of the most massive halos ($M_{\rm dm}>10^{7}\,\rm{M_{\odot}}$), with both the virial radius (given by Equation 1 below) and $\rho_{200/3}$ contour marked. Note that the $\rho_{200/3}$ density contour generally lies well within the virial radius for these halos. From these slices alone, it becomes apparent that a significant fraction of metals lie beyond the boundaries of halos, regardless of how they are defined.
@@
The top panel of Figure 3 shows the fraction of metal mass in halos $f_{\rm met}$, defining the boundary as either the virial radius or the $\rho_{200/3}$ density contour. Up until roughly 260 Myr ($z\approx 15.3$), $f_{\rm met}=0$%–5% within the $\rho_{200/3}$ contour, or $f_{\rm met}\approx 8$% on average within the virial radius, emphasizing that most of the metals produced are expelled out of the halo, and more importantly, out of star forming regions. It is not until after 260 Myr that metals begin to reaccrete to, or are retained within, the most massive halos following SN explosions. Even then, the metal fraction only reaches $f_{\rm met}\sim$31%–35%, however, this increase is rapid and may reach even higher fractions in a short amount of time following the end of the simulation. Of the metals in all halos, $92\%$ reside in the four most massive halos in the simulation. For comparison, $55\%$ of gas in star-forming halos resides in the four most massive halos, although only $0.8\%$ of gas resides in any halo in the box.
@@ 4 Metal Loss from Halos
The explosions of the first SNe quickly expel gas and metals from halos too small to retain their material. The fraction of metals retained by or, conversely, lost from each halo and the efficiency of mixing of the remaining metals determines when and how quickly star formation from enriched gas takes place in that halo or elsewhere.
Figure 4 shows that, at $z\lesssim 20$, the first instances of Pop III star formation occur in halos with masses of $M_{\rm dm}\sim 10^{5}$$10^{6}\,\rm M_{\odot}$, while, at $z\gtrsim 20$, the mass range narrows somewhat to $\sim 2$$5\times 10^{5}\,\rm M_{\odot}$. These results are consistent with those from Correa Magnus et al. (2024), which finds that the first minihalos that form Pop III stars are $M_{\rm dm}\sim 10^{5}\,\rm M_{\odot}$, and following nearby radiative feedback, the mass threshold for first Pop III star formation increases to $M_{\rm dm}\sim 10^{6}\,\rm M_{\odot}$. We also compare this finding with the model for critical halo mass $M_{\rm crit}$ for Pop III formation with $v_{bc}=0$ and $J_{LW}\sim 0-1\,J_{21}$ from Figure 3 of Kulkarni et al. (2021). The Kulkarni et al. (2021) model predicts a similar threshold mass to that demonstrated in Aeos, as well as a scatter of 0.2 dex in halo masses that is also shown in Figure 4 at lower redshifts. Given the short lifetimes of Pop III stars, Figure 4 suggests that the first Pop III SNe exploded in low-mass minihalos with $M_{\rm dm}<10^{6}\,\rm M_{\odot}$.
@@
The metal loss fraction of each individual parent halo (halos that are not part of a more massive halo) relative to the halo mass is shown in Figure 5111An animation of this plot over time is published online alongside this paper. The animated figure shows the metal loss fraction (top) and gas mass (bottom) evolution of individual halos as they grow in dark matter mass. As time progresses, more halos with SNe appear on the plot. In the top plot, halos with $M_{\rm dm}\lesssim 10^{7}\rm M_{\odot}$ generally initially appear with values of $\ell<1$, but quickly lose metals until $\ell\approx 1$. Halos with $M_{\rm dm}\gtrsim 10^{7}\rm M_{\odot}$ show progressively decreasing $\ell$. In the bottom plot, halos without SNe cluster around the cosmic baryon fraction with some scatter, continuously increasing in $M_{\rm dm}$ with time. Halos with SNe generally appear on the plot with gas mass significantly lower than the cosmic baryon fraction. As time progresses and halos grow in $M_{\rm dm}$, halos that have had SNe slowly increase in $M_{g}$ until halos $M_{\rm dm}\gtrsim 3\times 10^{6}$ have gas masses within an order of magnitude of the cosmic baryon fraction., color-coded by the number of SN that have exploded in the halo by the final redshift, $z=14.3$. Most halos have lost $\sim$100% of their injected metals, which appears to be independent of the number of SNe that had previously exploded in that halo. Those halos that have not lost all metals can be divided into two groups:
@@ 4.1.1 Characteristic Evolution
Figure 6 shows the time evolution in dark matter, baryon fraction $f_{b}$, metal mass, and metallicity respectively within the virial radius of the most massive progenitor of these four halos. Black circles represent instances of new SNe, with sizes scaled based on the number of new SNe. Magenta points indicate merger events.
All four halos show a similar dark matter mass evolution, originating very early on, within 50 Myr of each other and maintaining masses within 1 order of magnitude. Despite their similar dark matter evolution, they have widely varying gas evolution, with baryon fractions at times differing by up to 5 orders of magnitude. Rapid early decreases in $f_{b}$} are associated with the explosion of the first SNe.
Halos 1 and 2, which reach the lowest baryon fractions, have among the lowest dark matter masses at the minimum baryon fraction, hinting that the retention of material is tied to the depth of the potential well of the halo, which is closely tied to $T_{\rm vir}$ and is redshift dependent. At their $f_{b}$ minima, Halos 1 and 2 have $T_{\rm vir}\approx 798$ K ($v_{c}\sim 4.7\,\rm km\,s^{-1}$) and $1056$ K ($v_{c}\sim 5.4\,\rm km\,s^{-1}$) respectively, indicating shallower potentials than Halos 0 and 3, which have $T_{\rm vir}\approx 3295$ ($v_{c}\sim 9.5\,\rm km\,s^{-1}$) and $2800$ K ($v_{c}\sim 8.8\,\rm km\,s^{-1}$) respectively. Although Halo 0 has a similar dark matter mass to the other halos at the time of its first SN, it is in a stage of rapid dark matter growth, which slows the loss of gas and brings in new gas, keeping the baryon fraction from dropping more than an order of magnitude.
The time evolution of metal mass also demonstrates interestingly similar characteristics between these four halos. All halos, at the time of their first SNe, show a rapid spike in their metal mass, immediately followed by a metal loss of 2–3 orders of magnitude. Similar characteristics can be seen in the metallicity evolution. Subsequent SNe over the next $\sim$ 25 Myr also result in the loss of metal mass, after which halos begin to steadily acquire metals and gas from accretion, SNe, and winds. Most importantly, SN explosions after this point no longer appear to drive metal mass out of the halo, although they also no longer appear to contribute greatly to the total metal mass of the halo. This may indicate that, while the bulk of the metal mass, which is being accreted, is not being expelled from the halo, SNe yields are still escaping the halo. This is supported by the flat (or in the case of Halo 2, decreasing) metallicity with time, indicating that the ratios of metals to gas are not increasing as would be expected if yields injected by SNe were fully retained. As a result, metallicities only begin to increase near the end of the simulation.